OpenStudy (anonymous):

[(30 + 6) − 32] ÷ 9 • 2

OpenStudy (anonymous):

2/9

OpenStudy (anonymous):

if the dot means decimal point, then 10/23

OpenStudy (anonymous):

30+6-32=4 4/9.2=10/23

OpenStudy (chaise):

[(30 + 6) − 32] ÷ 9 • 2 (1/9)x[(30+6)-32 x 2 (1/9)x[36-32] x 2 (1/9)x[4] x 2 (1/9) x 8 8/9